Campbell Manufacturing Company uses two departments to make its products. Department I is a cutting department that is machine intensive and uses very few employees. Machines cut and form parts and then place the finished parts on a conveyor belt that carries them to Department II, where they are assembled into finished goods. The assembly department is labor intensive and requires many workers to assemble parts into finished goods. The company’s manufacturing facility incurs two significant overhead costs: employee fringe benefits and utility costs. The annual costs of fringe benefits are $264,000 and utility costs are $192,000. The typical consumption patterns for the two departments are as follows:

Department I Department II Total
Machine hours used 15,700 4,300 20,000
Direct labor hours used 5,300 10,700 16,000


The supervisor of each department receives a bonus based on how well the department controls costs. The company’s current policy requires using a single allocation base (machine hours or labor hours) to allocate the total overhead cost of $456,000.

Assume that you are the plant manager and have the authority to change the company’s overhead allocation policy. Formulate an overhead allocation policy that would be fair to the supervisors of both Department I and Department II. Compute the overhead allocations for each department using your policy.

Costs Department I Department II
Fringe benefits
Utility
Total $0 $0
Answer #1

The company must used activity based costing to allocate the overhead costs.

Cost per machine hour for the utility costs = 192,000 / 20,000 =9.6 per hour

Cost per labor hour for fringe benefits = 264,000 / 16,000 = 16.5 per hour

Costs Department I Department II
Fringe benefits 150720 41280
[15700*9.6] [4300*9.6]
Utility 87450 176550
[5300*16.5] [10700*16.5]
Total 238170 217830